# EntroPE: Entropy-Guided Dynamic Patch Encoder for Time Series Forecasting

### v1

@@ Abstract
Transformer-based models have significantly advanced time series forecasting, with patch-based input strategies offering efficiency and improved long-horizon modeling. Yet, existing approaches rely on temporally-agnostic patch construction, where arbitrary starting positions and fixed lengths fracture temporal coherence by splitting natural transitions across boundaries. This naive segmentation often disrupts short-term dependencies and weakens representation learning. In response, we propose EntroPE (Entropy-Guided Dynamic Patch Encoder), a novel, temporally informed framework that dynamically detects transition points via conditional entropy and dynamically places patch boundaries. This preserves temporal structure while retaining the computational benefits of patching. EntroPE consists of two key modules, namely an Entropy-based Dynamic Patcher (EDP) that applies information-theoretic criteria to locate natural temporal shifts and determine patch boundaries, and an Adaptive Patch Encoder (APE) that employs pooling and cross-attention to capture intra-patch dependencies and produce fixed-size latent representations. These embeddings are then processed by a global transformer to model inter-patch dynamics. Experiments across long-term forecasting benchmarks demonstrate that EntroPE improves both accuracy and efficiency, establishing entropy-guided dynamic patching as a promising new paradigm for time series modeling. Code is available at https://github.com/Sachithx/EntroPE.
@@ 1 Introduction
Time series analysis is fundamental to numerous critical applications, including electricity load forecasting (Gasparin et al., 2022), financial market prediction (Fischer & Krauss, 2018), healthcare monitoring (Reis & Mandl, 2003), and environmental surveillance (Smith, 1989). However, accurately modeling time series data remains challenging due to inherent noise, complex temporal dependencies, and irregular patterns spanning multiple time horizons (Lim & Zohren, 2021; Torres et al., 2021). The transformer architecture (Vaswani et al., 2017) has revolutionized sequence modeling across domains through self-attention mechanisms, leading to significant advances in time series forecasting. Despite their promise in learning long-term temporal relationships, time series transformers remain suboptimal due to unique complexities such as non-stationarity, seasonality, varying temporal granularities (Wen et al., 2023; Zeng et al., 2023).
@@
Recent advances in time series transformers include Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022), which address long-sequence modeling challenges through sparse attention mechanisms, decomposition techniques, and improved positional encodings while operating on point-wise inputs. Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) introduced a patch-based input representation, dividing raw sequences into fixed-length patches and achieving substantial improvements in computational efficiency and long-term forecasting performance. This demonstrated the critical impact of input tokenization strategies on transformer performance in time series. However, subsequent approaches have predominantly adopted similar temporally-agnostic patching schemes, which may not fully capture the temporal coherence and statistical properties inherent in time series data.
@@
Such temporally-agnostic patching introduces critical technical challenges. Fixed-length patching creates inconsistent input representations between training and inference phases. During training, time series samples from different time periods result in patches that begin at various temporal positions relative to underlying patterns (e.g., a patch might start mid-trend in one sample but at a trend beginning in another). However, during inference, patches are extracted from predetermined positions in the input sequence, creating a systematic distribution shift where the model encounters patch configurations it rarely or never saw during training. Further, predetermined patch boundaries arbitrarily segment coherent temporal structures without regard for underlying dynamics (see Fig. 1(a)). For example, a gradual trend change or seasonal transition may be split across multiple patches, breaking the natural temporal dependencies that are crucial for accurate pattern recognition and forecasting.
Figure 1: (a) Temporally-agnostic, and (b) Temporally-informed (proposed) patching with calculated entropies for ETTh1 sample.
@@
These challenges directly impact model performance through two mechanisms. Fragmented temporal patterns lead to incomplete intra-patch representations, where semantically related time points are separated across different patches and processed independently. Additionally, the train-inference mismatch reduces the model’s ability to generalize, as it must extrapolate to patch configurations with different statistical properties than those encountered during training. This is particularly detrimental for capturing rapid transitions and fine-grained temporal dynamics that require consistent temporal context.
@@
Unlike prior work that treats patching as a static preprocessing step, we propose EntroPE, an Entropy-Guided Dynamic Patch Encoder, which integrates dynamic boundary detection as a core architectural component. Our approach employs entropy-based criteria to identify natural temporal transition points, ensuring patch boundaries align with the underlying temporal structure rather than arbitrary positions (see Figure 1(b)). We then employ adaptive encoding mechanisms to process these variable-length patches while preserving intra-patch dependencies. This temporally-informed approach addresses both the train-inference mismatch and boundary fragmentation issues inherent in temporally-agnostic patching methods.
The main contributions of this paper are summarized as follows:
  • We introduce the first information-theoretic approach to patching for time series forecasting, where conditional entropy is used to place boundaries that respect temporal causality and predictive difficulty.
  • We design an adaptive encoder that converts variable-length patches into fixed representations via pooling and cross-attention, enabling efficient batch computation.
@@
  • We demonstrate consistent gains across diverse benchmarks, showing that entropy-guided dynamic patching is a practical and generalizable direction for time series transformers.
@@ 2 Related Work
Time Series Transformers and Long-Term Forecasting. The transformer architecture has fundamentally transformed sequence modeling across domains, with time series forecasting being no exception. Early adaptations like Informer (Zhou et al., 2021) addressed the quadratic complexity of self-attention in long sequences through ProbSparse attention mechanisms, while Autoformer (Wu et al., 2021) and FEDformer (Zhou et al., 2022) introduced decomposition-based approaches and frequency domain modeling to capture complex temporal dependencies. These foundational works established transformers as viable architectures for time series tasks but operated on point-wise inputs, leading to computational challenges and a limited ability to capture local temporal patterns efficiently. The fundamental challenge these approaches faced was the trade-off between sequence length and computational tractability. Point-wise processing, while preserving fine-grained temporal information, resulted in prohibitively long input sequences for practical applications, particularly in long-term forecasting scenarios where historical contexts spanning hundreds or thousands of time steps are crucial for accurate predictions.
@@
Patch-Based Input Strategies. Inspired by vision transformers (Dosovitskiy et al., 2021), patch-based input strategies have been widely adopted in time series forecasting. Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) segment time series into fixed-length patches that serve as input tokens. This design achieves a quadratic reduction in sequence length while preserving local semantic information through channel-independent processing, demonstrating that tokenization strategies can significantly improve both computational efficiency and forecasting accuracy. Building on this foundation, Crossformer (Zhang & Yan, 2023) introduced hierarchical mechanisms tailored for modeling both intra-patch and inter-patch dependencies. CARD (Xue et al., 2024) extended patching to multivariate forecasting, while patch-independence approaches (Lee et al., 2024) revealed superior representations in self-supervised settings which further highlights the importance of patch-based input strategies for time series modeling. Beyond Transformer architectures, xPatch (Stitsyuk & Choi, 2025) combined dual-flow MLP–CNN designs with exponential decomposition, whereas Pathformer (Chen et al., 2024) and PatchMLP (Tang & Zhang, 2025) leveraged hierarchical processing to capture multi-scale relationships.
However, fixed-length patching suffers from important limitations. Uniform segmentation can fragment coherent patterns such as seasonal transitions or trend shifts. During training, windows are sampled by shifting step by step so all possible boundaries are eventually seen, whereas inference fixes a single segmentation that may not align with natural transitions, weakening representation learning and harming generalization. In response, MSPatch (Cao et al., 2025) introduces multi-scale patch segmentation to capture both short- and long-term dependencies, mitigating the fragmentation of coherent temporal patterns. HDMixer (Huang et al., 2024) proposes a Length-Extendable Patcher (LEP) module that adjusts patch lengths via bi-linear interpolation to enrich boundary information, though it remains constrained by predetermined patch counts and may introduce interpolation-induced uncertainty. Beyond patch resizing, foundation models have explored alternative input constructions such as MOIRAI (Woo et al., 2024) leverages frequency-based patch sizing to preserve periodic structures, while Chronos (Ansari et al., 2024) bypasses patching entirely by treating individual time points as tokens, albeit at a substantially higher computational cost.
While recent works have advanced patch construction strategies, significant gaps remain. Existing adaptive methods often overlook causal structure and temporal coherence. This prevents them from capturing optimal boundaries guided by temporal uncertainty and predictive difficulty. Given the highly dynamic nature of time series across domains, a principled and fully adaptive approach is required. Moreover, efficient training capabilities are required, particularly due to the variable number of patches in each sample. In contrast, dynamic patching schemes in other domains, such as natural language processing, have shown superior performance and efficiency. For instance, the Byte Latent Transformer (Pagnoni et al., 2025) employes byte level inputs instead of traditional word tokenization, and dynamically decide patches as a part of the architeture. Motivated by these advances, our work introduces entropy-guided dynamic boundary detection that respects temporal causality and predictive uncertainty, coupled with adaptive encoding mechanisms specifically designed to preserve temporal dependencies within variable-length patches.
3 Methodology
3.1 Problem Formulation
Given a multivariate time series $X=[x_{1},x_{2},\ldots,x_{C}]\in\mathbb{R}^{C\times L}$ with $C$ channels and look-back length $L$, the goal is to predict $\hat{Y}\in\mathbb{R}^{C\times T}$ over forecast horizon $T$. Following the channel-independence principle from PatchTST (Nie et al., 2023), we deliberately disregard cross-channel dependencies and apply a shared function $f_{\phi}:\mathbb{R}^{L}\to\mathbb{R}^{T}$ (with shared parameters $\phi$) to each channel independently: $\hat{\mathbf{y}}_{c}=f_{\phi}(x_{c}),\quad c=1,2,\ldots,C$.
This design choice isolates the temporal modeling capabilities from cross-variate relationships, allowing us to focus purely on the effectiveness of our dynamic patching mechanism within each univariate series. The multivariate prediction is constructed by stacking channel-wise results:
$\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{C}]\in\mathbb{R}^{C\times T}$
@@ 3.1 Problem Formulation
Figure 2: Comprehensive architecture of EntroPE. The model processes input through: (A) Entropy-Based Dynamic Patcher - A small causal transformer calculates entropy at each time point to identify boundaries where predictive uncertainty is high; (B) Adaptive Patch Encoder - Cross-attention layers aggregate intra-patch dependencies into fixed-size global embeddings; (C) Fusion Decoder - Cross-attention combines global patch context with local encoder hidden states for accurate forecasting.
3.2 Forward Process
The forward process of EntroPE (see Fig. 2) transforms raw multivariate time series into forecasts through a sequence of interconnected components. The raw input multivariate time series $X\in\mathbb{R}^{B\times C\times L}$, where $B$, $C$, and $L$ represent batch size, number of channels (variables), and sequence length respectively, is first processed following the channel independence principle, converting it to $X^{ci}\in\mathbb{R}^{(B\cdot C)\times L}$ for independent processing of each variable. As the next step, to mitigate the statistical difference between training and testing distributions, we integrate the Reversible Instance Normalization layer (Kim et al., 2021) which outputs $X^{Norm.}\in\mathbb{R}^{(B\cdot C)\times L}$.
Following the Chronos tokenization approach (Ansari et al., 2024) (Appendix A.6), we discretize the continuous input sequence into uniform bins based on a fixed-size vocabulary $V$, producing $X^{tok}$. The tokenized sequence $X^{tok}$ is then processed by the EDP (explained in Sec. 3.3), which generates a patch mask $M\in\mathbb{R}^{B\cdot C\times L}$ indicating optimal patch boundary locations. Using this patch mask, the original continuous input $X^{ci}$ is segmented into variable-length patches. These patches are fed into the APE (explained in Sec. 3.4), which generates patch embeddings while aggregating intra-patch dependencies.
The encoder produces two key outputs, namely, (1) patch embeddings $P^{in}\in\mathbb{R}^{B\cdot C\times\hat{p}\times d_{p}}$, which serve as input embeddings for the global transformer, and (2) encoder hidden states $H^{enc}\in\mathbb{R}^{B\cdot C\times L\times d_{t}}$ that capture temporal information at the original sequence resolution for subsequent decoding head. Here, $\hat{p}$ is not fixed due to the dynamic nature of the patching, $d_{t}$ and $d_{p}$ are time point-wise embedding and patch embedding dimensions, respectively.
The patch embeddings $P^{in}$ are processed by the Global Transformer, which learns inter-patch long-term dependencies and outputs enriched patch embeddings $P^{out}\in\mathbb{R}^{B\cdot C\times\hat{p}\times d_{p}}$. Finally, the Fusion Decoder (FD) combines the global transformer outputs ($P^{out}$) with the encoder hidden states ($H^{enc}$) through cross-attention, flattens the enriched representations, and projects them through a linear head and Instance De-Normalization layer to generate predictions for the future $T$ time steps, producing the final forecast $\hat{Y}\in\mathbb{R}^{B\times C\times T}$.
3.3 Entropy-Based Dynamic Patcher (EDP)
As shown in Fig. 2(A), our dynamic patch boundary detection leverages a lightweight causal transformer to identify temporal transitions based on predictive uncertainty. The process operates in two independent phases, namely model pre-training for temporal pattern learning and entropy-guided patch boundary detection with frozen weights.
Entropy Model Pre-training.

### v2

@@ Abstract
Patch-based transformers have emerged as efficient and improved long-horizon modeling architectures for time series modeling. Yet, existing approaches rely on temporally-agnostic patch construction, where arbitrary starting positions and fixed lengths fracture temporal coherence by splitting natural transitions across boundaries. This naive segmentation often disrupts short-term dependencies and weakens representation learning. We propose a novel Entropy-Guided Dynamic Patch Encoder (EntroPE), as a temporally informed framework that dynamically detects transition points via conditional entropy and dynamically places patch boundaries. This preserves temporal structure while retaining the computational benefits of patching. EntroPE consists of two key modules, namely an Entropy-based Dynamic Patcher (EDP) that applies information-theoretic criteria to locate natural temporal shifts and determine patch boundaries, and an Adaptive Patch Encoder (APE) that employs pooling and cross-attention to capture intra-patch dependencies and produce fixed-size latent representations. Extensive experiments on long-term forecasting, classification, and anomaly detection demonstrate that the proposed method improves both accuracy and efficiency, establishing entropy-guided dynamic patching as a promising new paradigm for time series modeling.
@@ 1 Introduction
Figure 1: Comparison of patching strategies. (a) Temporally-agnostic fixed-length patching places boundaries at arbitrary positions, fragmenting coherent temporal patterns (highlighted regions show split transitions). (b) Our temporally-informed approach places boundaries at entropy peaks, where conditional entropy scale indicates natural transition points, preserving temporal coherence within patches.
@@
Time series analysis is fundamental to numerous critical applications, including electricity load forecasting (Gasparin et al., 2022), financial market prediction (Fischer and Krauss, 2018), healthcare monitoring (Reis and Mandl, 2003), and environmental surveillance (Smith, 1989). However, accurately modeling time series data remains challenging due to inherent noise, complex temporal dependencies, and irregular patterns spanning multiple time horizons (Lim and Zohren, 2021; Torres et al., 2021). The transformer architecture (Vaswani et al., 2017) has revolutionized language and sequence modeling across domains through self-attention mechanisms, leading to significant advances in time series analysis. Despite their promise in learning long-term temporal relationships, time series transformers remain suboptimal due to unique complexities such as non-stationarity, seasonality, varying temporal granularities (Wen et al., 2023; Zeng et al., 2023).
@@
Recent advances in time series transformers include Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022), which address long-sequence modeling challenges through sparse attention mechanisms, decomposition techniques, and improved positional encodings while operating on point-wise inputs. Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) introduced a patch-based input representation, dividing raw sequences into fixed-length patches and achieving substantial improvements in computational efficiency and performance. This demonstrated the critical impact of input tokenization strategies on transformer performance in time series. However, subsequent approaches have predominantly adopted similar temporally-agnostic patching schemes, which may not fully capture the temporal coherence and statistical properties inherent in time series data.
@@
These patching strategies treat time as temporally-homogeneous, partitioning sequences at arbitrary positions without regard for underlying temporal structure. Temporally-agnostic patching introduces two fundamental challenges. They create inconsistent input representations between training and inference phases. While training observes patches starting at diverse temporal offsets, result in patches that begin at various temporal positions relative to underlying patterns (e.g., a patch might start mid-trend in one sample but at a trend beginning in another). However, during inference, patches are extracted from predetermined positions in the input sequence, creating a systematic distribution shift in patch configurations. Further, predetermined patch boundaries arbitrarily segment coherent temporal structures without regard for frequent underlying dynamics (see Fig. 1(a)). For example, a frequently occurring gradual trend change or seasonal transition may be split across multiple patches, breaking the natural temporal dependencies that are crucial for accurate pattern recognition.
@@
These challenges directly impact model performance. Fragmented temporal patterns lead to incomplete intra-patch representations, where semantically related time points are separated across different patches and processed independently. Additionally, the train-inference mismatch reduces the model’s ability to generalize, as it must extrapolate to patch configurations with different statistical properties than those encountered during training. This is particularly detrimental for capturing rapid transitions and fine-grained temporal dynamics that require consistent temporal context.
@@
We argue that patch boundaries should align with points of elevated predictive uncertainty, where the past provides insufficient information to determine the future. Formally, such moments are characterized by high conditional entropy of the next observation given historical context. Segmenting time at these entropy peaks localizes regime changes at patch boundaries while yielding patches whose interiors remain predictively coherent. Motivated by this principle, we introduce EntroPE, an Entropy-Guided Dynamic Patch Encoder, which integrates uncertainty-aware boundary detection as a core architectural component. Unlike prior approaches that fix patch length or rely on heuristic change-point detectors, EntroPE derives segmentation from information-theoretic predictive structure, ensuring patch boundaries align with natural temporal transitions (Fig. 1(b)). We then employ adaptive encoding mechanisms to process these variable-length patches while preserving intra-patch dependencies. This temporally-informed approach addresses both the train-inference mismatch and boundary fragmentation issues inherent in temporally-agnostic patching methods.
The main contributions of this paper are:
  • We propose EntroPE, the first information-theoretic framework for dynamic patching in time series transformers.
  • We introduce an adaptive segmentation mechanism that balances modeling accuracy and attention complexity by dynamically controlling patch resolution.
@@
  • We demonstrate consistent improvements across forecasting, classification, and anomaly detection, establishing entropy-guided dynamic patching as a practical and generalizable paradigm for time series analysis.
Code is publicly available at https://github.com/Sachithx/EntroPE.
@@ 2 Related Work
Transformers have significantly advanced time series modeling by enabling long-range dependency learning through self-attention. Early adaptations such as Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), and FEDformer (Zhou et al., 2022) improved scalability via sparse attention, decomposition, and frequency-domain modeling. However, these approaches operate on point-wise tokens, resulting in prohibitively long sequences and a fundamental trade-off between temporal resolution and computational tractability.
@@
Patch-Based Input Representations. Inspired by Vision Transformers (Dosovitskiy et al., 2021), patch-based methods reduce sequence length while preserving local temporal semantics. Triformer (Cirstea et al., 2022) and PatchTST (Nie et al., 2023) pioneered fixed-length temporal patching, demonstrating substantial gains in efficiency and forecasting accuracy. Subsequent work extended this paradigm through hierarchical dependency modeling (Crossformer (Zhang and Yan, 2023)), multivariate forecasting adaptation (CARD (Xue et al., 2024)), and patch-independent self-supervised learning (Lee et al., 2024). Beyond transformers, xPatch (Stitsyuk and Choi, 2025), Pathformer (Chen et al., 2024), and PatchMLP (Tang and Zhang, 2025) explore CNN- and MLP-based hierarchical architectures to capture multi-scale temporal patterns.
Adaptive and Multi-Granularity Patching. Recent efforts relax fixed patch sizes through adaptive or multi-scale segmentation. MSPatch (Cao et al., 2025) employs multi-resolution patches, HDMixer (Huang et al., 2024) enables length-extendable interpolation, and MOIRAI (Woo et al., 2024) determines patch size in the frequency domain. AdaPatch (Liu et al., 2025a) mitigates non-stationarity via patch-level normalization and reconstruction, but operates over fixed segmentation rather than learning boundary placement. APN (Liu et al., 2025c) introduces time-aware adaptive patching for irregular time series through soft window aggregation, emphasizing data-density regularization instead of causality driven segmentation. SRSNet (Wu et al., 2025) constructs a selective representation space by scoring, selecting, and reassembling candidate patches, but operates over dense sliding-window patch pools rather than discovering causally meaningful boundary locations. TimeMosaic (Ding et al., 2025) further adapts patch granularity based on temporal heterogeneity and motif reuse, optimizing representation efficiency and decoding specialization through segment-wise prompting. While these methods improve flexibility, their segmentation criteria remain largely heuristic or representation-driven, prioritizing encoding density, frequency structure, or compression objectives over the underlying predictive or causal structure of the time series.
Causal and Information-Theoretic Perspectives on Segmentation. A complementary line of research argues that segmentation should reflect changes in predictive structure. In NLP, Byte Latent Transformer (Pagnoni et al., 2025) demonstrates that architecture-integrated dynamic tokenization improves efficiency by adapting token boundaries to information content. However, existing time series methods have not systematically grounded patch boundary selection in information-theoretic measures of predictive uncertainty or temporal causality.
In contrast to prior work that optimizes patching for representational efficiency or frequency structure, EntroPE places patch boundaries at peaks of conditional predictive entropy. This identifies moments where the past becomes insufficient to predict the future, typically corresponding to regime shifts or structural transitions. This principled segmentation preserves temporal coherence within patches, localizes uncertainty at boundaries, and eliminates train-inference patch mismatch. By coupling entropy-guided dynamic patching with adaptive encoding for variable-length segments, our method introduces a fundamentally new criterion for patch construction based on predictive uncertainty rather than heuristic granularity.
3 Method
3.1 Preliminaries: Problem Formulation
Given a multivariate time series $X=[x^{(1)},\ldots,x^{(C)}]\in\mathbb{R}^{C\times L}$ with $C$ channels and look-back length $L$, we address three fundamental tasks, namely forecasting, classification, and anomaly detection. Following the channel-independence principle (Nie et al., 2023), our architecture $f_{\phi}$ processes each channel independently to obtain channel-wise representations $\mathbf{z}^{(c)}=f_{\phi}(x^{(c)})$ for $c=1,\ldots,C$. Here, $f_{\phi}$ encompasses the complete processing pipeline: entropy-guided dynamic patching, adaptive patch encoding, global transformer, and fusion decoder (detailed in Sec. 3.3-3.6), producing representations $Z=[\mathbf{z}^{(1)},\ldots,\mathbf{z}^{(C)}]\in\mathbb{R}^{C\times d}$ where $d$ is the final representation dimension. $Z$ is projected via linear layer based on specific task.
@@ 3.1 Preliminaries: Problem Formulation
Figure 2: Comprehensive architecture of EntroPE. The model processes input through: (A) Entropy-Based Dynamic Patcher - A small causal transformer calculates entropy at each time point to identify boundaries where predictive uncertainty is high; (B) Adaptive Patch Encoder - Cross-attention layers aggregate intra-patch dependencies into fixed-size global embeddings; (C) Fusion Decoder - Cross-attention combines global patch context with local encoder hidden states for accurate predictions.
3.2 Overall Architecture
Our approach consists of two distinct phases (Fig. 2), (1) entropy model pre-training for uncertainty quantification, (2) task-specific training with entropy guided dynamic patching. Crucially, the entropy model parameters remain frozen during task-specific training, serving solely to provide patch boundaries.
3.3 Entropy Model Pre-training
As shown in Fig. 2(A), our dynamic patch boundary detection leverages a lightweight causal transformer to identify temporal transitions based on predictive uncertainty. Given a time series dataset $\mathcal{D}_{\text{pretrain}}=\{X^{(i)}\}_{i=1}^{N}$ where $X^{(i)}\in\mathbb{R}^{C\times L}$, we first quantize each continuous sequence into discrete tokens. For a univariate sequence $x=[x_{1},\ldots,x_{L}]$, we apply quantization $q:\mathbb{R}\to\mathcal{V}$ (defined in Eq. 38) to obtain tokenized sequence $\tau=[\tau_{1},\ldots,\tau_{L}]$ where $\tau_{t}=q(x_{t})\in\mathcal{V}$ and $|\mathcal{V}|=V$.
A critical distinction. Unlike Chronos (Ansari et al., 2024), which operates entirely in the quantized domain and suffers from range limitations during dequantization, we use quantization only for computing entropies and input projection. The downstream prediction heads project the output-embeddings, not quantized tokens. This decoupling eliminates the quantization range constraint as our predictions are not limited to the discrete vocabulary. Quantization serves purely as a preprocessing step for the entropy model (Sec. 3.3), while the actual time series transformer and prediction heads operate on continuous representations throughout.
Determining Quantization Range. To avoid manual threshold selection, we determine bin boundaries $\{b_{1},\ldots,b_{V-1}\}$ from training data statistics. Let $X_{\text{train}}$ denote z-score normalized training values. We compute empirical quantiles:
$q_{\text{low}}=Q_{\epsilon/2}(X_{\text{train}}),\quad q_{\text{high}}=Q_{1-\epsilon/2}(X_{\text{train}})$ (1)
where $\epsilon=0.005$ captures 99.5% of the distribution. The symmetric quantization radius is $R=\max(|q_{\text{low}}|,|q_{\text{high}}|)$, and the interval [R,R][-R,R] is uniformly divided into $V$ bins.